OpenStudy (anonymous):

Hmm, not sure how to do this seeing as how they are multiple questions... gimme a sec

OpenStudy (anonymous):

Simplify \[\sqrt{4x ^{2}+28x+49}\]

OpenStudy (anonymous):

Simplify \[\sqrt[4]{(-2x)^{4}}\]

OpenStudy (anonymous):

Add and Simplify \[\sqrt{12x+12} + \sqrt{3x+3}\]

OpenStudy (anonymous):

Subtract and Simplify \[5\sqrt{3y ^{3}} - \sqrt{12y}\]

OpenStudy (anonymous):

Multiply and Simplify \[\sqrt{18x} \times \sqrt{12x}\]

OpenStudy (anonymous):

Rationalize the Denominator and Simplify \[-4\sqrt{7} \div \sqrt{5} \times \sqrt{3}\] (THE DIVISION SYMBOL MEANS A FRACTION)

OpenStudy (anonymous):

Rationalize the Denominator and Simplify \[\sqrt{3} + 5\div 7+\sqrt{3}\] (THE DIVISION SYMBOL MEANS A FRACTION)

OpenStudy (anonymous):

Write with rational exponents and simplify \[\sqrt[4]{8x ^{3}y ^{2}}\]

OpenStudy (anonymous):

Solve \[\sqrt{5-3x} - 6=0\]

OpenStudy (anonymous):

Find the absolute value \[\left| -4+3i \right|\]

OpenStudy (anonymous):

I know these are a lot but I would really appreciate it if you could help me with as many as you can manage, I took this test after being absent for like two weeks and now I'm trying to catch up... these are all the problems I got wrong :( please help me, **I need to see your work so I can understand**

OpenStudy (istim):

Maybe someone could help you with one of the questions, and from that you derive how to solve the others.

OpenStudy (anonymous):

The problem with that is that they aren't all the same :/ I thought of that too but yeah as many as you could help with would be greatly appreciated

OpenStudy (istim):

For the top one, It's either a complex trinomial, or using the quadratic formula. There's only one answer though; the positive one.

OpenStudy (istim):

I did horrible in this class by the way.

OpenStudy (anonymous):

Lol, I got 2x + 9 \[\sqrt{7x}\] but it was marked wrong so idk

OpenStudy (istim):

I dunno.

OpenStudy (istim):

Sorry bro. I'm a lil rattled and also fuzzy on this subject (cause I didn't do enough work there).

OpenStudy (istim):

I think that some of this stuff is common knowledge though, just in a different setting.

OpenStudy (istim):

You good?

OpenStudy (anonymous):

Yeah, I'm like halfway done already lol thanks though

OpenStudy (istim):

Ok. Sorry fer disturbing you then (DON'T RESPOND).